OpenStudy (exceptionlearner):

Solve (x-2)4 (cube)=81

OpenStudy (anonymous):

\[(4(x-2))^3=81\] Is that what you are saying?

OpenStudy (exceptionlearner):

Yes

OpenStudy (anonymous):

Cube root both sides \[((4(x-2))^3)^{\frac{1}{3}}=(81)^{\frac{1}{3}}\]